2017-02-01 23:20 UTC: We are investigating the issue and will update the post as soon as we have further details.
2017-02-02 01:16 UTC: Web367 has suffered a catastrophic disk failure. We’ve recovered some data from the server, and have the remaining data in our backups. We’re presently restoring the machine to new hardware and will notify affected customers with the new server details.
2017-02-02 02:48 UTC: The file transfer is 50% complete.
2017-02-02 05:17 UTC: The file transfer is complete. We’re still working on restoring database passwords and permissions. If you notice a problem with your site, please open a support ticket for immediate assistance.
2017-02-02 15:00 UTC: We were able to recover all the data on Web367 and brought the original machine back online.
We are investigating the issue and will update the post as soon as we have further details.
2017-01-23 23:46 UTC: We’re running a filesystem check on Web340 at this time.
2017-01-24 00:40 UTC: The disk on Web340 is completely unrecoverable. We’re going to use the most recent backup to restore the machine to new hardware. We’ll email affected customers with the new server details.
2017-01-24 02:18 UTC: We’re transferring data from the backup server to the new hardware at this time.
2017-01-24 02:28 UTC: The new server details were just emailed to all affected customers.
2017-01-24 02:45 UTC: About 30% of the affected customers have been restored so far.
2017-01-24 05:37 UTC: About 62% of the affected customers have been restored so far.
2017-01-24 09:20 UTC: Most sites should now be back online. Large files are still being restored
2017-01-24 12:00 UTC: Large files are still being restored.
2017-01-24 19:06 UTC: Large files are still being restored.
2017-01-24 23:15 UTC: Some very large files for a very small number of accounts are still being restored.
2017-01-25 01:58 UTC: All accounts on Web340 have been restored to the new server. If you’re having any problems with your site on the new hardware, then please open a support ticket and our team will assist you.
2016-12-15 18:26 UTC: The server is currently not responding. We are investigating the issue and will update this post as soon as we have further details.
2016-12-15 18:34 UTC: The front end nginx server is up and operational at this time. Normal service has been restored. Our System Administrators will be monitoring to make sure that it starts normally.